OpenStudy (anonymous):

@Loveiskey18 Checking account A charges a monthly service fee of $20 and a wire transfer fee of $3, while checking account B charges a monthly service fee of $30 and a wire transfer fee of $2. How many transfers would a person have to have for the two accounts to cost the same? A. 21 B. 10 C. 0 D. 31

OpenStudy (anonymous):

If we are being realistic here, the answer is obviously 0. However this is math: c = 20 + 3t c = 30 + 2t 20 + 3t = 30 + 2t Solve for t: 20 + 3t - 2t = 30 + 2t - 2t 20 + t = 30 20 - 20 + t = 30 - 20 t = 10